OpenStudy (jdanmath):

Binomial distribution: n=19, p=0.07, q= 0.93 x=2 or more how do I find the probability of x being 2 or more?

OpenStudy (mathmate):

@jdanmath Do you know the equation for the probability for a particular value of x?

OpenStudy (jdanmath):

yes but that would mean i have to find for each x value

OpenStudy (jdanmath):

and then add them up. is there an easier way?

OpenStudy (mathmate):

Yes, you can do that, find P(x), x=2,3,4,5,6,7,8,....19. BUT...

OpenStudy (mathmate):

do you know the value of \(\sum_{i=0}^{n} P(i)\)?

OpenStudy (jdanmath):

no

OpenStudy (mathmate):

It means the sum of probabilities for i=0 to i=19 (in this case). Give it some thought?

OpenStudy (jdanmath):

i see okay

OpenStudy (mathmate):

So what is the sum?

OpenStudy (jdanmath):

190

OpenStudy (mathmate):

@jdanmath Probabilities can never exceed 1, nor less than zero. In this case, i can only take on values between 0 and 19 (inclusive), so the sum of the probabilities is 1, which means certainty. Based on that, you can simplify you calculations using the fact: P(0)+P(1)+.... +P(19)=1 so P(2)+P(3)+P(4)+....+P(19)=1-(P(0)+P(1)) To solve the problem, you can either do the left-hand side, or the right-hand side, using the formula I gave above for binomial distributions. Make your choice and finish the calculations! :)

OpenStudy (jdanmath):

got the right answer this time.. thank you so much!

OpenStudy (mathmate):

You're welcome! :)